# Embodying Physical Computing into Soft Robots

### v1

@@ 4 Algorithmic Physical Computing
Most importantly (and most relevant to this paper), the bistable mechanism can serve as a mechanical analog to transistors, functioning as one-bit memory units by switching between two stable configurations. These configurations can be mapped to binary states ‘0’ and ‘1’ based on the input force, pressure, or displacement, enabling the construction of logic gates and sequential logic circuits using entirely mechanical components  (?, ?, ?, ?, ?). Therefore, mechanical bi-stability provide a fundamental means to encode, store, and process information within a robot’s physical body. Typically, these robotic algorithmic computers directly borrow the design and architecture of CMOS-based systems, but they can be inherently energy efficient, retaining their state without continuous power input.
@@
To better illustrate the working principle, we detail a fluidic and algorithmic computer based on a reprogrammable metamaterial processor (Fig. 4a). The processor comprises identical bistable unit cells whose elastomeric chambers snap at defined pressure thresholds, converting vacuum and atmospheric pressure input into binary states (?). As a result, a bistable unit cell with a clever tubing design can function like a resistor, enabling the construction of complex logical circuitry. In the example in Fig. 4(a), 24 unit cells are connected into a soft processor consisting of two SR latches, a 2–4 demultiplexer, and four ring oscillators (each linked to a soft robotic finger). The soft processor and the fingers are all powered by one constant vacuum pressure.
@@
The entire system can reversibly switch between four different operation modes, each of which corresponds to the oscillatory bending of one finger. The operator can choose the operating mode via manually pressing the input cells of the SR latches (input encoding). For example, if the first figure is activated initially, one can press the $S_{2}$’ cell to activate the second figure and then press the $R_{2}$’ cell to switch back. The outputs of the two SR latches are sent to the demultiplexer, so these 2 data lines are converted into 4. The resulting four outputs of the demultiplexer serve as the power source for the ring oscillators (programmed input-output evolution). Finally, the robotic fingers transform the output oscillatory pressures into mechanical bending (output decoding). Notably, the current operation mode persists even after the removal of the pressing force, owing to the ability of the SR latches to retain their logic states until updated by new inputs. Finally, one can re-arrange these fluidic unit cells to construct a new soft processor with different input-output mapping (re-programming).
@@
Figure 4: Physical Algorithmic Computing. (a) Reprogrammable metamaterial processors with robotic fingers: Fluidic unit-cells with 0-1 binary states are connected to create mechanical logic circuitry to control finger actions (image adapted from (?) with permission). (b) Complementary soft pneumatic valves: Piston-based, four-terminal modules are paired to achieve Boolean logic operation, non-volatile latches, and analog pressure regulation. Then they are integrated with sub-circuits to create ring oscillators and counters to control crawling robots and wearable devices (image adapted from  (?) with permission). (c) Soft-matter computer: Conductive-fluid receptors transduce spatiotemporal fluid patterns into electrical drives, which can realize analog filtering, amplification, and logic gates with simple composition. As a result, such conductive fluidic mechanism enables on-body control for locomotion, reflexive grasping, and behavior switching (image adapted from (?) with permission).
@@
In addition to the example above, there are several other attempts to copy the electronic logic circuitry into the fluidic domain and construct fluid Boolean gates with soft complementary valves, ring oscillators, and modular cells (?, ?, ?, ?, ?) (Fig. 4b). Besides the pressurized fluidics, algorithmic computing can also be implemented with other novel materials and multi-stable mechanisms. Here, we list four additional approaches: 1) Conductive fluidics: Conductive fluidic receptors (CFRs) embedded in soft structures can act as hybrid mechanical-electrical logic units, enabling soft matter computers to perform sensing, logic, and actuation all in one continuous system (?, ?, ?) (Fig. 4c). 2) Magnetic fluidics: Magnetic liquid metal droplets can create flexible and reconfigurable logic gates with decoupled input/output channels and multi-modal control using phase-state transitions (?). 3) Heat responsive materials: Mechanical logic has also been achieved using mechanical and multiplexed switches that integrate bistable beams and thermally responsive materials to perform logic operations and mechanical memory storage (?, ?, ?, ?). 4) Multi-stable mechanisms: Finally, algorithmic computing is also possible with pure elastic force and deformation. Unique architectures, such as counter-snapping metamaterials, provide logic behavior via geometric nonlinearity, where structural instability enables programmable stiffness transitions and collective switching sequences, making them useful for timing and computation (?). Recent advances in modular chiral origami metamaterials further expand this logic repertoire by introducing multi-stable and reprogrammable architectures that can store information through mechanically encoded hysteresis and non-commutative state transitions (?).
@@ 5 Perspective for Future Advancement
Since this perspective lies at the intersection of physical computing and soft robots, it is intuitive to ask questions about the future direction using the “supply-and-demand” analogy. On the supply side: “Are there any newly available capabilities in physical computing that can be deployed for soft robotics?” On demand side: “What additional computing power would future soft robots require?” By surveying the current physical computing studies, one can discover many unique approaches that could be integrated into soft robots in the future (examples in Fig. 5). Here, we highlight three promising topics.
Figure 5: Future directions for embodying physical computing in soft robots: One can continue to advance the computing kernel’s density and capacity in soft robots by adopting new strategies to process encoded inputs, either locally or through a centralized kernel. To illustrate some of the promising concepts, we envision an octopus-inspired soft robot with computational capabilities distributed across its tentacles as well as in its brain, following the mechanical computing framework described in this manuscript. Each component can encode, process, and decode data (with mechanical memory reserved for storage). Starting from the top right tentacle and moving clockwise: Bistable soft shells enable rule-changeable logic operations (image adapted from (?) with permission); Information processing during transmission via non-dispersive mechanical solitary waves (image adapted from (?) with permission); A mechanical neural network offloads computation and can be attached to the robot’s skin (image adapted from (?) with permission); Mechanical analog-to-digital converters can be embedded in the tentacles (image adapted from (?) with permission). In the robot’s brain, miniaturized physical circuits mimic an algorithmic logic unit (ALU) (image adapted from (?) with permission). Finally, reprogrammable and non-volatile mechanical memories can store data either with magnetic (left, image adapted from (?) with permission) or thermal principles (right, image adapted from (?) with permission).
@@
An advanced and autonomous robot should also be able to memorize the operator’s instructions or the knowledge from its interaction with the working environment. In this regard, we have seen some promising examples of memory in the mechanical metamaterial domain. That is, by combining responsive materials (memory encoding/decoding) and elastic bistability (storage), mechanical metamaterials can achieve information storage via mechanical bits (m-bits), similar to their digital counterparts. These m-bits can be bistable origami structures (?) or tiles of bistable Kirigami units (?), forming 2D and 3D storage arrays. Examples include temperature-responsive bistable Kirigami units (?), which sequentially retrieve stored information in 3D arrays, and magnetic-responsive bistable elastic shell units (?), which enable on-demand re-programmability for 2D arrays. These configurations function as non-volatile mechanical memories and could be used in soft robotics in the future. A key challenge will be designing an integrative approach to encode information from the physical memory into the computing kernel, and subsequently decode and store the computing output into the physical memory device.
@@ 6 Conclusion: A Robust Route to Mechanical Intelligence
All these studies point to an exciting route to mechanical intelligence, where a soft robot can achieve intelligent behaviors, such as information perception, memory, and decision making, in the mechanical domain with minimal electronics. With a more advanced physical computing kernel, a soft robot might be able to acquire information from the surrounding environment, memorize it, and decide on the action plan, all in a highly integrated manner. This vision of soft robots with embodied physical computers hinges upon the ongoing convergence of many engineering disciplines, like mechanical metamaterials, computing theory, advanced manufacturing, and interdisciplinary design.

### v2

@@ 4 Algorithmic Physical Computing and Mechano-Logic
Most importantly (and most relevant to this paper), the bistable mechanism can serve as a mechanical analog to transistors, functioning as one-bit memory units by switching between two stable configurations. These configurations can be mapped to binary states ‘0’ and ‘1’ based on the input force, pressure, or displacement, enabling the construction of logic gates and sequential logic circuits using entirely mechanical components  (?, ?, ?, ?, ?). Therefore, mechanical bi-stability provides a fundamental means to encode, store, and process information within a robot’s physical body. Typically, these robotic algorithmic computers directly borrow the design and architecture of CMOS-based systems, but they can be inherently energy efficient, retaining their state without continuous power input.
@@
To better illustrate the working principle, we detail a fluidic and algorithmic computer based on a reprogrammable metamaterial processor (Fig. 4a). The processor comprises identical bistable unit cells whose elastomeric chambers snap at defined pressure thresholds, converting vacuum and atmospheric pressure input into binary states (?). As a result, a bistable unit cell with a clever tubing design can function like a resistor, enabling the construction of complex logical circuitry. More specifically
  • Kernel: In the example, the kernel has 24 unit cells that are connected into a soft processor, including two SR latches, a 2–4 demultiplexer, and four ring oscillators (each linked to a soft robotic finger). The soft processor and the fingers are all powered by one constant vacuum pressure. The entire system can reversibly switch between four different operation modes, each of which corresponds to the oscillatory bending of one finger.
@@
  • Input encoding: The operator can choose the operating mode via manually pressing the input cells of the SR latches (input encoding). For example, if the first figure is activated initially, one can press the $S_{2}$’ cell to activate the second figure and then press the $R_{2}$’ cell to switch back. The outputs of the two SR latches are sent to the demultiplexer, so these 2 data lines are converted into 4. The resulting four outputs of the demultiplexer serve as the power source for the ring oscillators.
  • Output decoding: The robotic fingers transform the output oscillatory pressures into mechanical bending. Notably, the current operation mode persists even after the removal of the pressing force, owing to the ability of the SR latches to retain their logic states until updated by new inputs.
  • Re-programming: Finally, one can re-arrange these fluidic unit cells to construct a new soft processor with different input-output mapping.
@@
Figure 4: Physical Algorithmic Computing. (a) Reprogrammable metamaterial processors with robotic fingers: Fluidic unit-cells with 0-1 binary states are connected to create mechanical logic circuitry to control finger actions (adapted from (?), $C$ BY 4.0). (b) Complementary soft pneumatic valves: Piston-based, four-terminal modules are paired to achieve Boolean logic operation, non-volatile latches, and analog pressure regulation. Then they are integrated with sub-circuits to create ring oscillators and counters to control crawling robots and wearable devices (adapted from (?), $C$ BY-NC-ND 4.0). (c) Soft-matter computer: Conductive-fluid receptors transduce spatiotemporal fluid patterns into electrical drives, which can realize analog filtering, amplification, and logic gates with simple composition. As a result, such conductive fluidic mechanism enables on-body control for locomotion, reflexive grasping, and behavior switching (image adapted from (?) with permission). Note that all scale bars are approximate.
@@
In addition to the example above, there are several other attempts to copy the electronic logic circuitry into the fluidic domain and construct fluid Boolean gates with soft complementary valves, ring oscillators, and modular cells (?, ?, ?, ?, ?, ?, ?) (Fig. 4b). Besides the pressurized fluidics, algorithmic computing can also be implemented with other novel materials and multi-stable mechanisms. Here, we list four additional approaches: 1) Conductive fluidics: Conductive fluidic receptors (CFRs) embedded in soft structures can act as hybrid mechanical-electrical logic units, enabling soft matter computers to perform sensing, logic, and actuation all in one continuous system (?, ?, ?) (Fig. 4c). 2) Magnetic fluidics: Magnetic liquid metal droplets can create flexible and reconfigurable logic gates with decoupled input/output channels and multi-modal control using phase-state transitions (?). 3) Heat responsive materials: Mechanical logic has also been achieved using mechanical and multiplexed switches that integrate bistable beams and thermally responsive materials to perform logic operations and mechanical memory storage (?, ?, ?, ?). 4) Multi-stable mechanisms: Finally, algorithmic computing is also possible with pure elastic force and deformation. Unique architectures, such as counter-snapping metamaterials, provide logic behavior via geometric nonlinearity, where structural instability enables programmable stiffness transitions and collective switching sequences, making them useful for timing and computation (?). Recent advances in modular chiral origami metamaterials further expand this logic repertoire by introducing multi-stable and reprogrammable architectures that can store information through mechanically encoded hysteresis and non-commutative state transitions (?).
@@
Challenges and opportunities of physical algorithmic computing: Compared to analog physical computers, algorithmic physical computing is quite versatile in that it can borrow many designs and working principles from well-established CMOS electronics. However, the reviewed examples above lag behind in terms of speed and scaling. Their computing speed is constrained by the relatively slow physical processes, such as pressurizing and venting of fluidic networks, deformation of thick elastomeric chambers, and, in some cases, heat diffusion through responsive materials. Using physical signals instead of electric ones also makes miniaturization more challenging: The finite size of multi-stable unit cells, the need for compliant interconnection devices, and the risk of mechanical crosstalk between unit cells make routing and isolation harder than in CMOS. Therefore, significant research efforts are necessary. For example, using advanced manufacturing technology can help minimize the unit cell size and enable more integrated packaging, thus speeding up physical computers (as we discuss later in Section 5). Regardless, physical algorithmic computing is still a desirable choice for small-scale logic and simple on-board control (as we discuss further in the conclusion section).
@@ 5 Perspective for Future Advancement
Since this perspective lies at the intersection of physical computing and soft robots, it is intuitive to ask questions about the future direction using the “supply-and-demand” analogy. On the supply side: “Are there any newly available capabilities in physical computing that can be deployed for soft robotics?” On demand side: “What additional computing power would future soft robots require?” Here, our unique perspective of dissecting soft robots into the encoding, kernel, and decoding layers can offer a systematic framework to introduce new computing concepts. For example, one can keep the encoding (e.g., sensor and input) and decoding components (e.g., actuator) the same, and “swap” the kernel with different designs that have new kinds of computing capacity. Alternatively, one can “upgrade” the kernel with a more advanced design. By surveying the current physical computing studies, one can discover many unique approaches that could be integrated into soft robots in the future (examples in Fig. 5).
Here, we highlight three most promising topics from three different angles: function, scale, and system integration. Regarding the function, the current physical computing has demonstrated an impressive ability to extract information from sensory signals and execute actuation commands. On the other hand, on-board non-volatile memory, a vital component in modern computing paradigms, has yet to be implemented in soft robots. Regarding scale, the current physical computing in soft robots is relatively large in terms of physical size. Minimizing their scale using high-precision manufacturing techniques might help advance the performance and reliability of physical computers to a new level, addressing some of the challenges in physical computing as we discussed earlier. Finally, regarding systematic integration, the current physical computing setup in soft robots primarily operates in a standalone manner. However, some integration with digital hardware (e.g., for long-distance communication) can enhance the overall capability. Therefore, a meaningful and integrated mechanical-electrical hybrid circuit can be advantageous. In the following section, we present recent studies in these three aspects.
Figure 5: Future directions for embodying physical computing in soft robots: One can continue to advance the computing kernel’s density and capacity in soft robots by adopting new strategies to process encoded inputs, either locally or through a centralized kernel. To illustrate some of the promising concepts, we envision an octopus-inspired soft robot with computational capabilities distributed across its tentacles as well as in its brain, following the mechanical computing framework described in this manuscript. Each component can encode, process, and decode data (with mechanical memory reserved for storage). Starting from the top right tentacle and moving clockwise: Bistable soft shells enable rule-changeable logic operations (image adapted from (?) $C$ BY 4.0); Information processing during transmission via non-dispersive mechanical solitary waves (image adapted from (?) $C$ BY 4.0); A mechanical neural network offloads computation and can be attached to the robot’s skin (image adapted from (?) with permission); Mechanical analog-to-digital converters can be embedded in the tentacles (image adapted from (?) with permission). In the robot’s brain, miniaturized physical circuits mimic an algorithmic logic unit (ALU) (image adapted from (?) $C$ BY 4.0). Finally, reprogrammable and non-volatile mechanical memories can store data either with magnetic (left, image adapted from (?) with permission) or thermal principles (right, image adapted from (?) $C$ BY 4.0). Note that all scale bars are approximate.
@@
An advanced and autonomous robot should also be able to memorize the operator’s instructions or the knowledge from its interaction with the working environment. In this regard, we have seen some promising examples of memory in the mechanical metamaterial domain. That is, by combining responsive materials (memory encoding/decoding) and elastic bistability (storage), mechanical metamaterials can achieve information storage via mechanical bits (m-bits), similar to their digital counterparts. These m-bits can be bistable elastic shell to realize mechano-fluidic memory (?), or bistable origami structures (?) or tiles of bistable Kirigami units (?), forming 2D and 3D storage arrays. Examples include temperature-responsive bistable Kirigami units (?, ?), which sequentially retrieve stored information in 3D arrays, and magnetic-responsive bistable elastic shell units (?), which enable on-demand re-programmability for 2D arrays. These configurations function as non-volatile mechanical memories and could be used in soft robotics in the future. A key challenge will be designing an integrative approach to encode information from the physical memory into the computing kernel, and subsequently decode and store the computing output into the physical memory device.
@@ 6 Conclusion: A Robust Route to Mechanical Intelligence
However, it is worth noting that all of the physically computing robots in this perspective remain at the proof-of-concept level. Implementing these exciting concepts into practical, real-world use still requires a significant amount of research efforts and system engineering.
Despite the rapid advances in this field, it is unlikely that physical computers embedded in soft robots can catch up with digital hardware in terms of computational speed and information density in the foreseeable future. Therefore, it is unrealistic to replace conventional digital hardware entirely with physical computing. Instead, engineers must answer the critical question: ”How much should we use physical computing?” ”Where to apply them?” and ”How can we seamlessly integrate physical computing with conventional digital hardware?”
For robots, physical computing is advantageous because of its softness, simplicity, and robustness. Therefore, it makes the most sense to use physical computing in the following three scenarios. 1) The targeted tasks are closely related to the robot’s physical body — this is why we have seen great success in locomotion generation and information extraction via direct physical interaction (using physical reservoir computing). 2) The robots need to be small and entirely soft — because physical computing could seamlessly integrate with the soft robotic body without the complexity of adding electronic components (e.g., using advanced 3D printing). 3) The working conditions are demanding — for example, fluidics-based computation is desirable for underwater operations, where electronics are vulnerable to damage. On the other hand, conventional digital electronics is more suitable for “over the distance” tasks, such as obstacle avoidance using vision data or long-distance communication with operators.
Therefore, the future of physically computing robots hinges upon two pillars: the continual advances in physical computing and its strategic integration with conventional digital hardware. As we discussed in Section 5, in the foreseeable future, we are likely to witness the creation of more powerful physical computing thanks to miniaturization and integrated memory capacity. With a more advanced physical computing kernel, a soft robot can acquire information from interacting with the surrounding environment, memorize the acquired knowledge, and execute the action plan, all in a highly integrated mechanical domain. On the other hand, new strategies will emerge to tightly integrate physical computing with digital computing via novel mechanical-electrical hybrid circuits, enabling physical computing robots to operate within large-scale automated systems. This vision of soft robots hinges on the ongoing convergence of various engineering disciplines, including mechanical metamaterials, computing theory, advanced manufacturing, and interdisciplinary design.